### v1

# MoReMouse: Monocular Reconstruction of Laboratory Mouse

Abstract
Laboratory mice play a crucial role in biomedical research, yet accurate 3D mouse surface motion reconstruction remains challenging due to their complex non-rigid geometric deformations and textureless appearance. Moreover, the absence of structured 3D datasets severely hinders the progress beyond sparse keypoint tracking. To narrow the gap, we present MoReMouse, the first monocular dense 3D reconstruction network tailored for laboratory mice. To achieve this goal, we highlight three key designs. First, we construct the first high-fidelity dense-view synthetic dataset for mice, by rendering our self-designed realistic Gaussian mouse avatar. Second, MoReMouse adopts a transformer-based feedforward architecture with triplane representation, achieving high-quality 3D surface generation from a single image. Third, we create geodesic-based continuous correspondence embeddings on mouse surface, which serve as strong semantic priors to improve reconstruction stability and surface consistency. Extensive quantitative and qualitative experiments demonstrate that MoReMouse significantly outperforms existing open-source methods in accuracy and robustness. Video results are available at https://zyyw-eric.github.io/MoreMouse-webpage/.
Figure 1: We present MoReMouse, the first monocular dense 3D reconstruction framework for laboratory mice. Given a single-view input (top-left, captured using an iPhone 15 Pro), MoReMouse predicts high-fidelity surface geometry and appearance using a transformer-based triplane architecture. The model is trained on a dataset constructed from real multi-view videos fitted to a Gaussian mouse avatar. Our method supports fast inference and outputs RGB renderings, semantic embeddings, and normal maps from a single image.
1 Introduction
Laboratory mice, as widely used model animals, are fundamental to disease modeling and neuroscience studies. Understanding their three-dimensional (3D) motion dynamics provides valuable insights for medical and life sciences[13].
Over the last decade, deep learning has revolutionized both 2D and 3D animal pose estimation, with methods such as DeepLabCut [24], SLEAP [29] and DANNCE [11] becoming standard tools across diverse species including mice. However, these methods are limited to sparse keypoint tracking and ignore important 3D surface motion. Although several papers have tried to build articulated mouse mesh models [6, 4, 1] and use them for mouse surface tracking [2, 17], they still face two key drawbacks. First, manually designed skinning weights are infeasible to represent the complex dynamic deformations of a mouse, resulting in severe self-penetration during its free movement. Second, existing mouse mesh fitting methods require heavy and sensitive multi-view optimization process, hindering its application for more convenient single view scenarios.
The rapid evolution of 3D generative AI has blurred the boundary between 3D reconstruction and 3D generation [39], where large scale 3D datasets such as Objaverse [10, 9] play a crucial role in this trend. Thanks to the datasets, Large Reconstruction Models (LRMs) [14, 39, 38] demonstrate single-view feedforward 3D reconstruction for general objects, significantly outperform traditional feedback approaches like DreamFusion [30] in both reconstruction quality and efficiency. Although LRMs have the potential to revolutionize monocular mouse reconstruction, they require ground truth 3D geometries or dense-view images for supervision, which is unavailable for laboratory mice. Additionally, mice exhibit highly dynamic non-rigid deformations, and their furry skins lack prominent visual features, making it more challenging to reconstruct surface motion of mice.
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse. MoReMouse is built upon transformer architectures and triplane representation, enabling efficient, feedforward 3D reconstruction. To compensate for the scarcity of dense-view 3D mouse dataset, we construct the first animatable Gaussian avatar of mouse using an articulated mouse model [6] and Gaussian Splatting [16]. Inspired by recent human avatar modeling [20], we propose a 2D UV-based intermediate mouse pose representation to drive Gaussian point clouds. Thanks for the free-view rendering ability and diverse pose space of the avatar, we build a synthetic 64-view dataset, enabling effective training of MoReMouse network.
Despite of the synthetic dataset, the framework of MoReMouse integrates multiple architectural and training enhancements inspired by LRMs. Specifically, we employ a two-stage training strategy, initially pretraining the model with triplane-NeRF [7, 25] representations, followed by fine-tuning with triplane-DMTet[35] for higher-fidelity reconstructions. To further improve the stability and robustness of dense surface reconstruction, we introduce geodesic-based feature embeddings, which enhance semantic consistency of surfaces across frames. In experiments, we compare MoReMouse with existing open-sourced feedforward large reconstruction models such as TripoSR [39], Triplane-GS [50], InstantMesh [45] and LGM [38] on our synthetic dataset, proving the superiority of MoReMouse for monocular mouse reconstruction and novel view synthesis. Qualitative results on real-world videos captured by hand-held IPhone further confirm the generalization ability of MoReMouse (Fig. 1). As a whole, MoReMouse represents a crucial step toward high-fidelity, efficient 3D reconstruction of mouse, and will facilitate broader applications in biomedical research. To summarize, our key contributions include:
  • We introduce MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse, surpassing existing open-source alternatives.
  • We develop the first Gaussian avatar model for mice and generate a large scale dense-view synthetic mouse dataset for training and evaluation.
  • We integrate geodesic-based continuous correspondence embeddings with feedforward 3D reconstruction, improving single-view reconstruction stability and quality.
2 Related work
2.1 3D Representation of Mouse
Previously, most 3D mouse motion capture toolboxes reconstruct mouse 3D keypoints by algebraic triangulation [27, 15, 12, 26], voxel-based triangulation [11] or even by markers [22, 23]. However, 3D keypoints provide only sparse geometric structures, neglecting dense surface information which is essential to represent highly dynamic mouse body shapes or interpret social touch information [2, 17]. To represent the mouse surface,  [5] proposes an articulated mesh model based on computed tomography scans. Although this model is useful for the motion capture of mouse surface from multiview videos [2], its manually designed skinning weight could not capture the realistic dynamic surface motion of a moving mouse. ArMo [4] presents a similar mouse mesh model, but lacks realistic geometric details. For the rat, which shares a shape similar to that of the mouse, researchers pay similar efforts to build articulated mesh model [1] yet face the same challenge. In this paper, we try to overcome the limitation of explicit articulated representation of mouse by reconstructing the implicit surface of it.
2.2 Monocular Reconstruction of Animals
Since the proposal of SMAL (Skinned Multi-Animal Linear) model which represents several quadrupedal species [52], predicting the pose and shape parameters of SMAL from images has dominated the monocular reconstruction of animals [53, 54, 3, 44, 32, 33, 21]. Although they have demonstrated satisfactory performance for species such as dogs, horses, cattle, and tigers, they cannot reconstruct rodents due to the limited shape space of SMAL. The AWOL model [51] further extends SMAL to represent rodents using CLIP [31] and supports monocular shape estimation. However, it still relies on a predefined articulated model, without the ability to express geometric deformation beyond linear blend skinning. Some recent papers try to build animal avatars from videos by augmenting the skinning weights or adding additional deformations. For example, BANMo [47] and RAC [48] build fully implicit cat and dog avatars using neural deformation field and differentiable rendering [25], Gart [18] and Animal Avatars [34] build dog avatars with SMAL geometric embedding and neural appearance. However, these methods all require a heavy optimization process during inference, and rely on dense pose estimation such as CSE [28] or SMAL parameters for initialization.
2.3 Large Reconstruction Model
Large reconstruction models have recently achieved substantial progress in generating high-quality 3D content [14, 19, 41, 46, 50, 49, 36, 37]. These models are typically trained in an end-to-end fashion to predict implicit 3D representations, often through triplane-based neural radiance fields. Some approaches operate from a single input image [14, 41, 46, 50, 49], while others use multi-view images generated by novel view synthesis techniques [19, 36]. Meanwhile, UNet-based architectures have also proven effective in generating multiplane or Gaussian representations [42, 38], providing alternative and efficient formats for 3D reconstruction.
However, these methods generally depend on access to 3D supervision, such as ground-truth shapes or high-quality multi-view data, during training. This reliance poses a significant challenge for in-the-wild applications, where diverse and scalable 3D datasets remain scarce compared to their 2D counterparts [8, 10, 43]. In addition, their reconstructions often exhibit limitations in geometric detail and texture quality, making them less suitable for scenarios that demand high visual fidelity.
3 Method
Our goal is to achieve dense geometry reconstruction of a mouse from a single viewpoint via a feed-forward 3D generative approach, where representation and supervision play crucial roles in ensuring high-quality generation.
To construct dataset for effective supervision (Sec. 3.1), we design the first Gaussian avatar for mouse (Sec. 3.1.2), which introduces an additional bias on top of the mesh model to better capture complex mouse surface deformations from videos (Sec. 3.1.3).
Our proposed MoReMouse model builds upon an LRM-like structure (Sec. 3.2) and exploits the strengths of both Triplane-NeRF for an initial coarse training phase and Triplane-DMTet for fine-grained training (Sec. 3.2.1). Furthermore, to enhance tracking stability and preserve surface consistency, we introduce a geodesic-distance-based feature embedding (Sec. 3.2.2). These embeddings serve as additional semantic priors, improving the robustness of dense reconstruction. The details of our training and optimization procedure are provided in (Sec. 3.2.3).
3.1 Synthetic Dataset using Gaussian Avatar
Figure 2: Visualization of the articulated mouse mesh and the learned geodesic embedding. Left: our mesh model consists of 140 articulated joints and 14,522 vertices, forming a kinematically consistent 3D structure for pose fitting. Right: color-coded visualization of the learned geodesic feature embedding, where principal components are mapped to HSV color space to enhance surface separability and promote reconstruction consistency.
3.1.1 Mouse Parametric Model
As shown in Fig 2, we employ the established mouse mesh model used in  [2] as avatar control signals. The model consists of $K=140$ articulated joints and 14,522 vertices, and conforms to Linear Blend Skinning (LBS) for pose control. The whole parameter set is denoted as $\Psi$, including joint rotations $\theta\in\mathbb{R}^{3K}$, global translation $t\in\mathbb{R}^{3}$, global rotation $r\in\mathbb{R}^{3}$, global scale $s\in\mathbb{R}$ and bone length deformation parameters $B\in\mathbb{R}^{21}$ (Please refer to  [2] for more details). We decompose $\Psi$ into local parameters $\Psi_{l}=\{\theta,B\}$ and global parameters $\Psi_{g}=\{r,t,s\}$ for the sake of simplicity.
3.1.2 Gaussian Mouse Avatar
Figure 3: Pipeline of the Gaussian Mouse Avatar Generation. The process begins with Mesh Fitting, followed by Gaussian model training.
Gaussian point clouds [16] use an explicit parameterization for 3D representation. Each Gaussian point is characterized by a mean position $\mu\in\mathbb{R}^{3}$ and an anisotropic covariance matrix $\Sigma$:
$G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}.$ (1)
To facilitate optimization, the covariance matrix is reformulated as $\Sigma=RSS^{T}R^{T}$, where $S$ is a diagonal scaling matrix controlled by a three-dimensional vector $s\in\mathbb{R}^{3}$, and the rotation matrix $R$ is represented by a normalized quaternion $q\in\mathbb{R}^{4}$. Additionally, each Gaussian encodes a color $c\in\mathbb{R}^{3}$ and an opacity $o\in\mathbb{R}$. During rendering, Gaussians within the pixel projection region are sorted by depth and composited using alpha blending.
Inspired by Animatable Gaussians [20], we employ texture-based deformation transfer to train a neural representation of the Gaussian point cloud avatar, incorporating bias correction beyond the skeletal model, as shown in Fig. 3. Our key idea is to parameterize the mouse UV texture, mapping each valid texel in the UV space to a corresponding Gaussian point, where the texture color represents the Gaussian’s position.
To obtain the canonical space representation, we set $\Psi_{g}=0$ and drive the mesh model using only $\Psi_{l}$, resulting in canonical vertex coordinates $V_{\Psi_{l}}$. The UV-mapped positions of these vertices are used to interpolate Gaussian locations in the UV texture via barycentric coordinates. This defines the baseline Gaussian position map $\mu_{\Psi_{l}}$:
$\mu_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}),$ (2)
where $\mu_{0}$ is the UV coordinate map in the neutral pose, and $\mathcal{D}_{\Psi_{l}}$ represents the LBS deformation mapping induced by the skeletal model.
During Gaussian optimization, we incorporate a CNN-based offset prediction module. A convolutional neural network StyleUNet [40] predicts the position offset map $\Delta\mu_{\Psi_{l}}$, conditioned on the UV coordinate map $\mu_{\Psi_{l}}$, to refine the Gaussian positions:
$\mu^{\prime}_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}+\Delta\mu_{\Psi_{l}}).$ (3)
The CNN also estimates additional Gaussian attributes, including color $c_{\Psi_{l}}$, opacity $o_{\Psi_{l}}$, covariance rotation parameters $q_{\Psi_{l}}$, and scaling parameters $s_{\Psi_{l}}$. Then the Gaussian Splatting rendering function generates the view-dependent rendered image $I_{c}^{\text{rend}}$:
$I_{c}^{\text{rend}}=GS(R_{c},T_{c},K_{c},\mu^{\prime}_{\Psi_{l}},c_{\Psi_{l}},q^{\prime}_{\Psi_{l}},s_{\Psi_{l}}),$ (4)
where $GS(\cdot)$ denotes the Gaussian splatting renderer, and $R_{c},T_{c},K_{c}$ are camera parameters.
3.1.3 Avatar Training
To train the Gaussian avatar, we define loss functions between the real image $I_{c}$ and the rendered image $I_{c}^{\text{rend}}$:
$\mathcal{L}$ $=\mathcal{L}_{1}(I_{c}-I_{c}^{\text{rend}})+\lambda_{\text{SSIM}}(1-\mathcal{L}_{\text{SSIM}}(I_{c},I_{c}^{\text{rend}}))$ (5)
$\quad+\lambda_{\text{LPIPS}}\mathcal{L}_{\text{LPIPS}}(I_{c},I_{c}^{\text{rend}}),$
where $\mathcal{L}_{\text{SSIM}}$ quantifies structural similarity, while $\mathcal{L}_{\text{LPIPS}}$ assesses perceptual similarity. The loss weights are set to $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{LPIPS}}=0.1$. Additionally, we enforce Total Variation (TV) loss to regularize Gaussian point cloud smoothness.
We use a 6-view mouse video sequence ”markerless_mouse_1” proposed in [11] for avatar training. Each sequence captures 18,000 frames of a single freely moving mouse in an open field at 1152x1024@100FPS, where the mouse mesh models are provided by  [2]. Note that, we only use 800 frames for training, which are uniformly sampled from the first 8000 frames of ”markerless_mouse_1”. Avatar training takes 400k steps.
3.2 MoReMouse Architecture
3.2.1 Transformer and Triplane Representation
Similar to LRM [14], MoReMouse adopts a Transformer and triplane-based architecture, designed for monocular 3D reconstruction. The model consists of three key components: an image encoder, an image-to-triplane decoder, and a triplane-based geometric representation that integrates both NeRF [25] and DMTet [35] for implicit and explicit geometry modeling.
First, the image encoder employs a pre-trained DINOv1 backbone to extract a latent feature vector from the input image. Then, a large Transformer decoder employs cross-attention to project these image features into a 3D triplane representation. This triplane-based geometric representation encodes color, density, and additional spatial attributes at any queried 3D location.
During the initial training phase, to regularize the convergence of the spatial density field, we do not introduce explicit geometry. Instead, we employ NeRF for differentiable volumetric rendering, allowing gradients to propagate throughout the entire space.
Once the density field has sufficiently converged, we introduce DMTet as an explicit geometric representation for further refinement. DMTet provides well-defined surfaces and enables direct geometric supervision.
The fine-scale structures of mice, particularly in the limbs and tail, present challenges for high-fidelity reconstruction in a feed-forward manner. To address this, we modify the LRM framework by increasing the triplane resolution to 64 to improve detail preservation. Additional architectural details are provided in the supplementary materials.
Figure 4: Architecture of the MoReMouse reconstruction network. Given a monocular input image, a DINO-V2 encoder extracts high-level image features, which are processed by a transformer-based decoder to generate a triplane representation. These triplanes, combined with learnable positional embeddings, are queried by a Multi-Head MLP to predict per-point color, density, and geodesic embedding. Rendering is performed via either NeRF (for volumetric supervision) or DMTet (for explicit geometry output), producing RGB images, color-coded embeddings, and surface meshes.
3.2.2 Feature Embedding for Surface Consistency
The reconstruction of mice poses significant challenges due to their black fur, which lacks distinctive surface features, and the presence of intricate fine-scale geometric details. Inspired by Animatable Gaussians [20], we incorporate feature embeddings into the mouse surface representation to enhance reconstruction stability. These embeddings serve as additional color channels during training, acting as supervisory signals for MoReMouse rendering.
Specifically, given the set of mouse model vertices $V\in\mathbb{R}^{13059\times 3}$, we compute the pairwise geodesic distances between vertices to obtain the geodesic distance matrix:
$G_{V}\in\mathbb{R}^{13059\times 13059},\quad G_{V}(i,j)=d_{\text{geo}}(V_{i},V_{j}),$ (6)
where $d_{\text{geo}}(V_{i},V_{j})$ represents the geodesic distance between vertices $V_{i}$ and $V_{j}$.
To learn a continuous correspondence embedding that captures the intrinsic structure of the mouse surface, we define the embedding matrix $E\in\mathbb{R}^{13059\times 3}$ and optimize it by minimizing the $L_{2}$ loss between the embedding-derived distance matrix $G_{E}$ and the geodesic distance matrix $G_{V}$:
$\mathcal{L}_{\text{geo}}=\sum_{i,j}\left\|G_{E}(i,j)-G_{V}(i,j)\right\|_{2}^{2},$ (7)
where $G_{E}(i,j)=\|E_{i}-E_{j}\|_{2}$ represents the Euclidean distance in the embedding space.
However, directly using these embeddings as color channels is suboptimal. The three-dimensional embeddings often form an elongated mouse-like shape in feature space, failing to fully utilize the available color space. Additionally, large dark or bright regions appear, lacking sufficient contrast against the black or white backgrounds used during rendering. To address this, we perform Principal Component Analysis (PCA) on the embeddings, selecting the first two principal components as the hue (H) and saturation (S) values in the HSV color space, while keeping the value (V) channel fixed at 1:
$H=\text{PCA}_{1}(E),\quad S=\text{PCA}_{2}(E),\quad V=1.$ (8)
This transformation improves the separability of the embeddings in the rendered images.
Notably, a simple rotation of the embeddings in color space can generate new feature mappings:
$E^{\prime}=R(\theta)E,$ (9)
where $R(\theta)$ is a rotation matrix. Extending the embeddings to higher dimensions could further improve reconstruction robustness. However, we leave the exploration of such extensions to future work.
3.2.3 Training and Optimization
Following the method outlined in Section 3.1.2, we generate freely renderable Mouse Avatars for each frame of the ”markerless_mouse_1” video sequence. As the dataset consists of 18,000 frames, we use the first 6,000 frames for training and the last 6,000 frames for testing. To construct the training set, we sample two sets of 64 viewpoints per frame, rendering 12,000 multi-view scenes in total. The optical axes of these viewpoints converge at the scene origin.
During training, two randomly selected viewpoints from each scene serve as the input-output pair. To mitigate spatial ambiguity, we fix the mouse position at the scene origin. In preprocessing, we also apply translation and scaling transformations to the input images, ensuring that the mouse centroid is aligned with the image center and normalized to a consistent scale. Experimental results show that MoReMouse generalizes well from training on fixed perspectives to varying viewpoints in real-world scenarios.
We employ a two-stage training strategy, where NeRF-based volumetric rendering is first trained for 60 epochs to ensure proper density field convergence, followed by DMTet-based explicit geometry modeling for an additional 100 epochs to refine surface details.
The total training loss is defined as:
$\mathcal{L}$ $=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{smoothL1}}\mathcal{L}_{\text{smoothL1}}$ (10)
$\quad+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{LPIPS}}\mathcal{L}_{\text{LPIPS}}.$
Here, $\mathcal{L}_{\text{MSE}}$ enforces pixel-wise supervision for both RGB values and feature embeddings, while $\mathcal{L}_{\text{depth}}$ ensures depth consistency within the valid object region. The smooth L1 loss $\mathcal{L}_{\text{smoothL1}}$ penalizes large RGB discrepancies, and $\mathcal{L}_{\text{mask}}$ applies binary cross-entropy supervision to the predicted opacity maps. The perceptual loss $\mathcal{L}_{\text{LPIPS}}$ measures high-level perceptual similarity in a deep feature space. The hyperparameters $\lambda$ control the relative importance of each term.
4 Experiments
4.1 Experimental Setup
We evaluate MoReMouse on two types of testing datasets: a synthetic benchmark and a real-world captured dataset.
Synthetic Evaluation Data.
We use the last 6,000 frames from the markerless_mouse_1 sequence to generate a synthetic multi-view dataset consisting of 4 orthogonal viewpoints. These views are rendered using our Gaussian avatar pipeline and serve as a controlled evaluation set where ground-truth camera parameters and rendering conditions are known.
Real Captured Data.
To assess performance under realistic conditions, we collect a real-world multi-view dataset using industrial cameras. This dataset contains 5,400 frames of freely moving mice captured from four calibrated viewpoints. The camera rotation angles relative to the reference view are $0^{\circ}$, $-33.39^{\circ}$, $-78.60^{\circ}$, and $-181.02^{\circ}$, respectively.
Baselines.
Due to the lack of monocular 3D reconstruction methods specifically designed for mice, we compare MoReMouse against four general-purpose single-view reconstruction baselines:
  • TripoSR [39]: A state-of-the-art open-source Large Reconstruction Model (LRM) known for strong single-view 3D generation performance.
  • Triplane-GS [50]: A LRM variant based on Gaussian representations.
  • LGM [38]: A multi-view diffusion-based method that reconstructs 3D Gaussian point clouds.
  • InstantMesh [45]: A mesh-based 3D reconstruction method using multi-view diffusion priors.
For synthetic data, we additionally evaluate the multi-view diffusion baselines (LGM and InstantMesh) using ground-truth views as input, which serves as an approximate upper bound for multi-view reconstruction.
Evaluation Protocol.
All methods are evaluated on the task of novel view synthesis, since obtaining ground-truth 3D geometry for mouse subjects is impractical. For each frame, a single view is used as input, and the remaining three views are used to evaluate the visual fidelity of synthesized images. We report three standard image-based metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS).
4.2 Experimental Results
Quantitative Results.
We report the evaluation results on the synthetic and real-world datasets in Table 1. For each metric, we highlight the top-3 performing methods using color shading, with darker tones indicating better performance.
Table 1: Quantitative results on both synthetic and real-world datasets. MoReMouse achieves the best performance across all metrics in both domains, significantly outperforming baseline methods in terms of reconstruction fidelity.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic Ours 22.027 0.9660 0.05279
TripoSR 13.673 0.8032 0.18255
Triplane-GS 18.049 0.9268 0.10151
LGM 14.460 0.8805 0.13274
InstantMesh 15.821 0.8987 0.11312
real Ours 18.422 0.9478 0.08674
TripoSR 11.518 0.8114 0.19672
Triplane-GS 16.789 0.9298 0.11002
LGM 15.215 0.9197 0.12838
InstantMesh 15.631 0.9175 0.11339
Across all three metrics for novel view synthesis, MoReMouse consistently outperforms the baseline methods, demonstrating its ability to effectively infer plausible mouse geometry and pose from a single image. Notably, in Table 2, MoReMouse uses only a single input view, while LGM and InstantMesh are evaluated with 4 and 6 views respectively. For LGM, the 4 input views happen to coincide with the 4 evaluation viewpoints, offering an idealized setting that may favor the method. Nonetheless, MoReMouse achieves comparable performance despite relying on significantly less input information, underscoring its strong generalization capabilities in monocular reconstruction scenarios.
We also observe that the quantitative metrics on the real-world dataset are generally lower than on the synthetic dataset, especially for PSNR. This is largely due to the camera-mouse configuration: in the real dataset, the cameras are fixed while the mouse moves, whereas MoReMouse and all baselines assume that the mouse is centered and aligned across input-output views. This mismatch introduces pixel-level alignment errors, reducing PSNR. However, the perceptual metrics SSIM and LPIPS remain high, indicating that MoReMouse still reconstructs plausible and correct mouse poses and shapes. This confirms that models trained on synthetic data can generalize to real-world mouse motion capture scenarios.
Table 2: Comparison with multi-view diffusion-based methods using ground-truth viewpoints on the synthetic dataset. Despite using only a single view, MoReMouse performs competitively with multi-view approaches that use 4 or 6 calibrated inputs.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours (single view) 22.027 0.9660 0.05279
LGM (4 views) 23.998 0.9698 0.06458
InstantMesh (6 views) 28.295 0.9831 0.02716
Qualitative Results.
To visually compare MoReMouse with baseline methods, we select three representative frames from both the synthetic and real-world sequences. As shown in Fig. 5 and Fig. 6, MoReMouse generates mice with coherent appearance and physically plausible postures, accurately capturing tail and limb movements. These results remain stable even though each frame is inferred independently, showing smooth motion continuity.
In contrast, baseline methods often exhibit structural distortions in novel views and inconsistent pose transitions across frames. Diffusion-based approaches such as InstantMesh and LGM tend to produce view-inconsistent geometry, likely due to the absence of strong priors on mouse anatomy. Single-view approaches like Triplane-GS and TripoSR frequently fail to reconstruct accurate poses, especially under extreme articulations or occlusions.
4.3 Ablation Study
4.3.1 Comparison with Finetuned TripoSR
In previous experiments, we compared MoReMouse with four general-purpose single-view 3D reconstruction baselines, demonstrating the effectiveness of our complete training pipeline. However, these reconstruction models were pre-trained on datasets that may not contain sufficient small animal or mouse-like samples, which could limit their performance on our mouse-specific benchmarks.
To assess whether the improvements of MoReMouse are due to architectural and representation innovations rather than only data augmentation, we fine-tune TripoSR on our mouse dataset using a similar training protocol. Specifically, we train the model for 120 epochs with a learning rate of $1\times 10^{-6}$. As shown in Table 3 and illustrated in Fig. 7, the fine-tuned TripoSR achieves improved performance compared to its original version, but still lags behind MoReMouse across most metrics. This suggests that our model architecture, including the triplane encoding and geodesic feature embedding, plays a critical role in achieving high-fidelity reconstructions.
Table 3: Comparison between MoReMouse and TripoSR finetuned on mouse data.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic Ours 22.027 0.9660 0.05279
TripoSR-Tuned 21.996 0.9676 0.06333
real Ours 18.422 0.9478 0.08674
TripoSR-Tuned 18.162 0.9461 0.09354
4.3.2 Effect of Feature Embedding
We evaluate the effectiveness of the geodesic feature embedding module by training a variant of MoReMouse without this component. As shown in Table 4 and visualized in Fig. 8, removing the embedding leads to noticeably degraded reconstruction quality, particularly around fine-scale structures.
Without embedding supervision, the model struggles to maintain surface consistency in anatomically delicate regions such as limbs and tails. As evident in the first, third, and fourth rows of Fig. 8, the full model produces sharper and more accurate reconstructions of the paws and extremities. In the second row, the embedded model also better preserves subtle texture details around the snout.These observations support our design choice to use surface-based semantic priors. The geodesic embedding encourages coherent geometry and finer detail recovery.
Figure 5: Qualitative comparison of novel view synthesis results on synthetic data. MoReMouse produces coherent and anatomically plausible reconstructions, accurately capturing mouse posture and geometry, while baseline methods often exhibit structural distortions.
Figure 6: Qualitative comparison of novel view synthesis results on real data. MoReMouse generalizes well to real-world mouse recordings, maintaining consistent geometry and appearance across views, while baseline methods often suffer from inconsistent shapes and misaligned postures.
Figure 7: Qualitative comparison between our MoReMouse method and a finetuned TripoSR baseline. MoReMouse produces sharper and more anatomically coherent results. In contrast, TripoSR—based on volumetric NeRF rendering—tends to generate blurry boundaries and over-smoothed geometry, especially around fine structures like the ears.
Figure 8: Qualitative comparison with and without geodesic feature embedding. The full model produces sharper and more consistent geometry, especially around limbs (rows 1, 3, 4) and snout texture (row 2). Embeddings help guide fine-grained surface details and improve anatomical accuracy.
Table 4: Ablation study on the effect of feature embedding. Removing embedding results in lower reconstruction quality, especially in perceptual similarity.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
synthetic full model 22.027 0.9660 0.05279
w/o embedding 21.805 0.9655 0.05467
real full model 18.422 0.9478 0.08674
w/o embedding 18.250 0.9469 0.08767
4.4 Failure cases
Figure 9 showcases typical failure cases on real-world data. The first row highlights the model’s inability to generalize to rare upright postures, due to the absence of such poses in the training set. This leads to implausible geometry and emphasizes the need for more pose-diverse avatar data.
In the second row, reconstruction fails when the tail occludes the head—especially under shaved-fur conditions—causing the model to misinterpret head structure. While our preprocessing removes background, we do not explicitly account for occlusions from scene objects or self-occlusion, which can degrade reconstruction fidelity.
Additionally, we observe that small regions such as feet and snouts are more prone to distortion, likely due to limited resolution in those areas. Future work may explore resolution-aware refinement or occlusion-robust modeling to address these limitations.
Figure 9: Representative failure cases. Top: failure due to lack of standing mouse data in training, leading to unrealistic reconstructions. Bottom: failure due to tail-induced occlusion from shaved fur, which interferes with head feature extraction.
5 Discussion
Summary
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model tailored for laboratory mice. Unlike existing single-view 3D methods designed for general objects or humans, MoReMouse addresses the unique challenges posed by small animals, including non-rigid deformation, low surface texture, and limited training data. Our approach integrates transformer-based triplane representations, a Gaussian avatar modeling pipeline, and geodesic surface feature embeddings to achieve robust and high-fidelity reconstructions from a single image. We demonstrate strong performance across both synthetic and real datasets, outperforming state-of-the-art reconstruction baselines in terms of visual quality and perceptual consistency. Ablation studies further validate the architectural contributions of our model beyond dataset-specific priors. We hope that MoReMouse can serve as a foundation for future research in computational ethology and non-rigid animal modeling, promoting more precise, efficient, and scalable behavioral analysis tools for biomedical research.
Limitations.
While MoReMouse offers a promising solution to monocular mouse reconstruction, several limitations remain that warrant future investigation. (1) Limited training diversity. Our training data is derived from a single video sequence of a single mouse under consistent lighting and appearance. This limits the generalization of the model to other mice with varying fur color, size, or under different lighting conditions. (2) Pose tracking in global coordinates. The current system assumes that the mouse is always centered in a canonical scene space. This design limits our ability to track full 3D poses in a real-world coordinate system.
Ethical Statement.
All procedures involving mice in this study were non-invasive and observational, involving only video recording of natural behaviors without any physical intervention or stress to the animals. The experiments were conducted in accordance with internationally recognized guidelines for laboratory animal care. Since the study involved only observational methods, ethical approval was not required by the relevant institutional animal care and local ethics authority. The mice were housed under standard conditions in compliance with local animal welfare regulations, and all efforts were made to minimize disturbance during filming. The study adhered to the 3Rs principle (Replacement, Reduction, Refinement) in animal research.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62125107) and the Shuimu Tsinghua Scholar Program (2024SM324).
Supplementary Material
1 MoReMouse Model Details
We summarize the architectural and training hyperparameters of the MoReMouse model in Table 5, covering key components such as the image encoder, triplane representation modules, multi-head MLP, NeRF and DMTet renderers, and training configurations.
Table 5: MoReMouse Architecture and Training Hyperparameters
Module Parameter Value
Image Tokenizer Type dinov2-base
Image resolution $378\times 378$
Feature channels $768$
Triplane Tokenizer Plane size $64\times 64$
Channels $512$
Backbone Input channels $512$
Attention layers $12$
Attention heads $16$
Attention head dimension $64$
Cross attention dimension $768$
Triplane Upsampler Input channels $512$
Output channels $80$
Output shape $3\times 80\times 128\times 128$
MultiHeadMLP Neurons $64$
Shared hidden layers $10$
Hidden layers for density/feature/deformation heads $1$
Renderer-NeRF Samples per ray $128$
Radius $0.87$
Density activation trunc_exp
Density bias $-1.0$
Renderer-DMTet SDF bias $-4.0$
Isosurface resolution $256^{3}$
Training Learning rate $1\text{e}{-5}$
Optimizer AdamW
LR scheduler CosineAnnealingLR
Warm-up steps $3000$
$\lambda_{\text{mse}}$, $\lambda_{\text{LPIPS}}$ $1.0$
$\lambda_{\text{mask}}$ $0.3$
$\lambda_{\text{smooth}}$, $\lambda_{\text{depth}}$ $0.2$

### v2

@@
 Abstract
-Laboratory mice play a crucial role in biomedical research, yet accurate 3D mouse surface motion reconstruction remains challenging due to their complex non-rigid geometric deformations and textureless appearance.
-Moreover, the absence of structured 3D datasets severely hinders the progress beyond sparse keypoint tracking.
-To narrow the gap, we present MoReMouse, the first monocular dense 3D reconstruction network tailored for laboratory mice.
-To achieve this goal, we highlight three key designs.
-First, we construct the first high-fidelity dense-view synthetic dataset for mice, by rendering our self-designed realistic Gaussian mouse avatar.
-Second, MoReMouse adopts a transformer-based feedforward architecture with triplane representation, achieving high-quality 3D surface generation from a single image.
-Third, we create geodesic-based continuous correspondence embeddings on mouse surface, which serve as strong semantic priors to improve reconstruction stability and surface consistency.
-Extensive quantitative and qualitative experiments demonstrate that MoReMouse significantly outperforms existing open-source methods in accuracy and robustness.
-Video results are available at https://zyyw-eric.github.io/MoreMouse-webpage/.
+Laboratory mice, particularly the C57BL/6 strain, are essential animal models in biomedical research.
+However, accurate 3D surface motion reconstruction of mice remains a significant challenge due to their complex non-rigid deformations, textureless fur-covered surfaces, and the lack of realistic 3D mesh models.
+Moreover, existing visual datasets for mice reconstruction only contain sparse viewpoints without 3D geometries.
+To fill the gap, we introduce MoReMouse, the first monocular dense 3D reconstruction network specifically designed for C57BL/6 mice.
+To achieve high-fidelity 3D reconstructions, we present three key innovations.
+First, we create the first high-fidelity, dense-view synthetic dataset for C57BL/6 mice by rendering a realistic, anatomically accurate Gaussian mouse avatar.
+Second, MoReMouse leverages a transformer-based feedforward architecture combined with triplane representation, enabling high-quality 3D surface generation from a single image, optimized for the intricacies of small animal morphology.
+Third, we propose geodesic-based continuous correspondence embeddings on the mouse surface, which serve as strong semantic priors, improving surface consistency and reconstruction stability, especially in highly dynamic regions like limbs and tail.
+Through extensive quantitative and qualitative evaluations, we demonstrate that MoReMouse significantly outperforms existing open-source methods in both accuracy and robustness.
+Project pagehttps://zyyw-eric.github.io/MoreMouse-webpage/
 Figure 1: We present MoReMouse, the first monocular dense 3D reconstruction framework for laboratory mice.
-Given a single-view input (top-left, captured using an iPhone 15 Pro), MoReMouse predicts high-fidelity surface geometry and appearance using a transformer-based triplane architecture.
-The model is trained on a dataset constructed from real multi-view videos fitted to a Gaussian mouse avatar.
-Our method supports fast inference and outputs RGB renderings, semantic embeddings, and normal maps from a single image.
+Given a single-view input (top-left, captured with an iPhone 15 Pro), MoReMouse predicts high-fidelity surface geometry and appearance within 0.9 seconds using a transformer-based triplane architecture (middle).
+To assist model training, we render a dense-view synthetic dataset by building the first Gaussian mouse avatar from sparse-view real videos (bottom-left).
+Our method outputs RGB renderings, semantic embeddings, and normal maps from a single image (right).
+Corresponding video results are provided in the supplementary material.
 1 Introduction
-Laboratory mice, as widely used model animals, are fundamental to disease modeling and neuroscience studies.
-Understanding their three-dimensional (3D) motion dynamics provides valuable insights for medical and life sciences[13].
-Over the last decade, deep learning has revolutionized both 2D and 3D animal pose estimation, with methods such as DeepLabCut [24], SLEAP [29] and DANNCE [11] becoming standard tools across diverse species including mice.
+Laboratory mice are widely used model organisms in biomedical research, especially for disease modeling, genetic studies, and neuroscience.
+Among them, the C57BL/6 strain is the most commonly used inbred line due to its well-characterized genome and relevance in neurodegenerative, oncological, and immunological research.
+Accurate 3D motion reconstruction of C57BL/6 mice is essential for behavioral phenotyping.
+Over the last decade, deep learning has revolutionized both 2D and 3D animal pose estimation, with methods such as DeepLabCut (mathis2018deeplabcut), SLEAP (pereira2022sleap) and DANNCE (dunn2021geometric) becoming standard tools across diverse species including mice.
 However, these methods are limited to sparse keypoint tracking and ignore important 3D surface motion.
-Although several papers have tried to build articulated mouse mesh models [6, 4, 1] and use them for mouse surface tracking [2, 17], they still face two key drawbacks.
-First, manually designed skinning weights are infeasible to represent the complex dynamic deformations of a mouse, resulting in severe self-penetration during its free movement.
-Second, existing mouse mesh fitting methods require heavy and sensitive multi-view optimization process, hindering its application for more convenient single view scenarios.
-The rapid evolution of 3D generative AI has blurred the boundary between 3D reconstruction and 3D generation [39], where large scale 3D datasets such as Objaverse [10, 9] play a crucial role in this trend.
-Thanks to the datasets, Large Reconstruction Models (LRMs) [14, 39, 38] demonstrate single-view feedforward 3D reconstruction for general objects, significantly outperform traditional feedback approaches like DreamFusion [30] in both reconstruction quality and efficiency.
-Although LRMs have the potential to revolutionize monocular mouse reconstruction, they require ground truth 3D geometries or dense-view images for supervision, which is unavailable for laboratory mice.
-Additionally, mice exhibit highly dynamic non-rigid deformations, and their furry skins lack prominent visual features, making it more challenging to reconstruct surface motion of mice.
-In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse.
-MoReMouse is built upon transformer architectures and triplane representation, enabling efficient, feedforward 3D reconstruction.
-To compensate for the scarcity of dense-view 3D mouse dataset, we construct the first animatable Gaussian avatar of mouse using an articulated mouse model [6] and Gaussian Splatting [16].
-Inspired by recent human avatar modeling [20], we propose a 2D UV-based intermediate mouse pose representation to drive Gaussian point clouds.
+To this end, template-based papers have tried to build articulated mouse mesh models (49_bolanos2021three) and used them for mouse 3D tracking (an2023three).
+However, the manually designed skinning weights are infeasible to represent the complex dynamic deformations of a mouse, resulting in severe self-penetration during its free movement.
+Moreover, mesh fitting methods require heavy and sensitive multi-view optimization, hindering its application for more convenient single view scenarios.
+Template-free animal reconstruction methods such as 3D-Fauna (li2024learning) lacks the knowledge of mouse-specific appearance and geometry, therefore has difficulty in recovering 3D C57BL/6 mice.
+The rapid evolution of 3D generative AI has blurred the boundary between 3D reconstruction and 3D generation, with large-scale datasets such as Objaverse (deitke2023objaverse) playing a crucial role in this trend.
+Fueled by such data, Large Reconstruction Models (LRMs) (hong2023lrm; tochilkin2024triposr; tang2024lgm) have demonstrated high efficiency and accuracy in single-view, feedforward 3D reconstruction of general objects.
+More recent generative frameworks, such as Trellis (xiang2025structured) based on rectified flow, and CLAY (zhang2024clay) based on diffusion, further push the scalability and controllability of 3D generation by learning powerful structured latent representations.
+Although these models show great promise, they do not perform well on laboratory mice due to the lack of large-scale 3D mice supervision or dense-view mice images.
+Moreover, as these methods only considers single-image cases, no temporally semantic correspondences are ensured for long term dynamic object reconstruction.
+In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model specifically designed for C57BL/6 laboratory mouse.
+MoReMouse is built upon transformer architectures and triplane representation, enabling efficient, feedforward 3D reconstruction, as shown in Fig.
+1.
+To compensate for the scarcity of dense-view 3D mouse dataset, we construct the first animatable Gaussian avatar of mouse (AGAM for short) using an articulated mouse model (49_bolanos2021three) and Gaussian Splatting (kerbl2023gaussian).
+Inspired by recent human avatar modeling (li2024animatable), we use a 2D UV-based intermediate mouse pose representation to drive Gaussian point clouds.
 Thanks for the free-view rendering ability and diverse pose space of the avatar, we build a synthetic 64-view dataset, enabling effective training of MoReMouse network.
-Despite of the synthetic dataset, the framework of MoReMouse integrates multiple architectural and training enhancements inspired by LRMs.
-Specifically, we employ a two-stage training strategy, initially pretraining the model with triplane-NeRF [7, 25] representations, followed by fine-tuning with triplane-DMTet[35] for higher-fidelity reconstructions.
-To further improve the stability and robustness of dense surface reconstruction, we introduce geodesic-based feature embeddings, which enhance semantic consistency of surfaces across frames.
-In experiments, we compare MoReMouse with existing open-sourced feedforward large reconstruction models such as TripoSR [39], Triplane-GS [50], InstantMesh [45] and LGM [38] on our synthetic dataset, proving the superiority of MoReMouse for monocular mouse reconstruction and novel view synthesis.
-Qualitative results on real-world videos captured by hand-held IPhone further confirm the generalization ability of MoReMouse (Fig.
-1).
-As a whole, MoReMouse represents a crucial step toward high-fidelity, efficient 3D reconstruction of mouse, and will facilitate broader applications in biomedical research.
+Based on the new synthetic dataset, the framework of MoReMouse integrates multiple architectural and training enhancements inspired by LRMs.
+Specifically, we employ a two-stage training strategy, initially pretraining the model with triplane-NeRF (chan2022efficient; mildenhall2021nerf) representations, followed by fine-tuning with triplane-DMTet(shen2021deep) for higher-fidelity reconstructions.
+To further improve the stability and robustness, we introduce geodesic-based feature embeddings, which enhance semantic consistency of surfaces across frames.
+We evaluate MoReMouse on both synthetic and real-world datasets.
+The synthetic benchmark is rendered using our AGAM, while the real-world evaluation uses a 4-view video sequence of single freely moving C57BL/6 mouse captured by PointGray industrial cameras.
+Results demonstrate the superior performance of MoReMouse in monocular mouse reconstruction and novel view synthesis.
+As a whole, MoReMouse represents a significant step toward practical and high-fidelity monocular 3D reconstruction of C57BL/6 mice, with potential to support broader research in computational modeling and analysis of small animals.
 To summarize, our key contributions include:
-  • We introduce MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse, surpassing existing open-source alternatives.
-  • We develop the first Gaussian avatar model for mice and generate a large scale dense-view synthetic mouse dataset for training and evaluation.
-  • We integrate geodesic-based continuous correspondence embeddings with feedforward 3D reconstruction, improving single-view reconstruction stability and quality.
+  • We introduce MoReMouse, the first monocular dense 3D reconstruction model tailored for laboratory mice, specifically the C57BL/6 strain, and show its superiority over existing open-source alternatives.
+  • We develop the first animatable Gaussian avatar model for mice and use it to generate a large-scale, dense-view synthetic dataset for training and evaluation.
+  • We integrate geodesic-based continuous correspondence embeddings into feedforward reconstruction, improving single-view surface stability and detail preservation.
 2 Related work
-2.1 3D Representation of Mouse
-Previously, most 3D mouse motion capture toolboxes reconstruct mouse 3D keypoints by algebraic triangulation [27, 15, 12, 26], voxel-based triangulation [11] or even by markers [22, 23].
-However, 3D keypoints provide only sparse geometric structures, neglecting dense surface information which is essential to represent highly dynamic mouse body shapes or interpret social touch information [2, 17].
-To represent the mouse surface,  [5] proposes an articulated mesh model based on computed tomography scans.
-Although this model is useful for the motion capture of mouse surface from multiview videos [2], its manually designed skinning weight could not capture the realistic dynamic surface motion of a moving mouse.
-ArMo [4] presents a similar mouse mesh model, but lacks realistic geometric details.
-For the rat, which shares a shape similar to that of the mouse, researchers pay similar efforts to build articulated mesh model [1] yet face the same challenge.
-In this paper, we try to overcome the limitation of explicit articulated representation of mouse by reconstructing the implicit surface of it.
-2.2 Monocular Reconstruction of Animals
-Since the proposal of SMAL (Skinned Multi-Animal Linear) model which represents several quadrupedal species [52], predicting the pose and shape parameters of SMAL from images has dominated the monocular reconstruction of animals [53, 54, 3, 44, 32, 33, 21].
-Although they have demonstrated satisfactory performance for species such as dogs, horses, cattle, and tigers, they cannot reconstruct rodents due to the limited shape space of SMAL.
-The AWOL model [51] further extends SMAL to represent rodents using CLIP [31] and supports monocular shape estimation.
-However, it still relies on a predefined articulated model, without the ability to express geometric deformation beyond linear blend skinning.
-Some recent papers try to build animal avatars from videos by augmenting the skinning weights or adding additional deformations.
-For example, BANMo [47] and RAC [48] build fully implicit cat and dog avatars using neural deformation field and differentiable rendering [25], Gart [18] and Animal Avatars [34] build dog avatars with SMAL geometric embedding and neural appearance.
-However, these methods all require a heavy optimization process during inference, and rely on dense pose estimation such as CSE [28] or SMAL parameters for initialization.
-2.3 Large Reconstruction Model
-Large reconstruction models have recently achieved substantial progress in generating high-quality 3D content [14, 19, 41, 46, 50, 49, 36, 37].
+1 3D Representation of Mouse
+Previously, most methods reconstruct mouse 3D keypoints by algebraic triangulation (35_nath2019using; 36_karashchuk2021anipose; han2024multi; monsees2022estimation), voxel-based triangulation (dunn2021geometric) or even by optical markers (marshall2021pair).
+However, 3D keypoints provide only sparse geometric structures, neglecting essential dense surface information.
+To represent the mouse surface,  (49_bolanos2021three) proposes an articulated mesh model based on computed tomography scans, and is used for multi-view mesh fitting (an2023three).
+ArMo (bohnslav2023armo) presents a similar mouse mesh model, but lacks true anatomical bones.
+For rats, which are similar to mice in shape, similar mesh models are built (aldarondo2024virtual; klibaite2025mapping).
+All these mesh models are limited to the manually designed skinning weights, which could not capture the realistic dynamic surface motion of mice.
+2 Monocular Reconstruction of Animals
+Since the proposal of SMAL model which represents several quadrupedal species (63_zuffi20173d), predicting the pose and shape parameters of SMAL from images has dominated the monocular reconstruction of animals (zuffi2019three; biggs2020wldo; xu2023animal3d; rueegg2022barc; lyu2025animer).
+Although they perform well for quadrupeds, they cannot reconstruct rodents due to the limited shape space of SMAL.
+The AWOL model (zuffi2024awol) extends SMAL to represent rodents using CLIP (radford2021learning), yet its pose and shape realism is still limited.
+Some recent papers try to build animal avatars from videos by augmenting the skinning weights or adding additional deformations, such as BANMo (yang2022banmo), Gart (lei2024gart) or AnimalAvatars (Sabathier2024AnimalAR).
+However, these methods all require a heavy optimization process during inference, and rely on dense pose estimation or accurate SMAL parameters for initialization.
+Template-free methods such as MagicPony (wu2023magicpony) and 3D Fauna (li2024learning) leverage collections of single-view images of deformable species captured in the wild to jointly learn canonical models and pose regressors.
+However, these approaches do not generalize well to mice due to domain and geometry gaps.
+3 Large Reconstruction Model
+Large reconstruction models have recently achieved substantial progress in generating high-quality 3D content (hong2023lrm; Li2023Instant3DFT; zou2024triplane; gslrm2024; sun2024dreamcraft3d++).
 These models are typically trained in an end-to-end fashion to predict implicit 3D representations, often through triplane-based neural radiance fields.
-Some approaches operate from a single input image [14, 41, 46, 50, 49], while others use multi-view images generated by novel view synthesis techniques [19, 36].
-Meanwhile, UNet-based architectures have also proven effective in generating multiplane or Gaussian representations [42, 38], providing alternative and efficient formats for 3D reconstruction.
-However, these methods generally depend on access to 3D supervision, such as ground-truth shapes or high-quality multi-view data, during training.
-This reliance poses a significant challenge for in-the-wild applications, where diverse and scalable 3D datasets remain scarce compared to their 2D counterparts [8, 10, 43].
-In addition, their reconstructions often exhibit limitations in geometric detail and texture quality, making them less suitable for scenarios that demand high visual fidelity.
+Recent methods improve quality and efficiency by learning generative models over compact latent codes derived via VAE-like encoders (hu2025simulating).
+Some focus purely on shape (zhang20233dshape2vecset; li2024craftsman3d), while others incorporate appearance (gupta20233dgen; xiong2024octfusion).
+CLAY (zhang2024clay) presents a unified system for geometry and PBR texture generation.
+Trellis (xiang2025structured) introduces structured latent codes via rectified flow transformers, jointly modeling geometry and appearance for versatile decoding into NeRFs, Gaussians, and meshes.
+However, these methods generally depend on access to 3D supervision (deitke2023objaverse; wu2023omniobject3d), such as ground-truth shapes or high-quality multi-view data, during training.
+The data scarcity is one of the main challenges for training an LRM for laboratory mice.
 3 Method
-Our goal is to achieve dense geometry reconstruction of a mouse from a single viewpoint via a feed-forward 3D generative approach, where representation and supervision play crucial roles in ensuring high-quality generation.
-To construct dataset for effective supervision (Sec.
-3.1), we design the first Gaussian avatar for mouse (Sec.
-3.1.2), which introduces an additional bias on top of the mesh model to better capture complex mouse surface deformations from videos (Sec.
-3.1.3).
-Our proposed MoReMouse model builds upon an LRM-like structure (Sec.
-3.2) and exploits the strengths of both Triplane-NeRF for an initial coarse training phase and Triplane-DMTet for fine-grained training (Sec.
-3.2.1).
-Furthermore, to enhance tracking stability and preserve surface consistency, we introduce a geodesic-distance-based feature embedding (Sec.
-3.2.2).
-These embeddings serve as additional semantic priors, improving the robustness of dense reconstruction.
-The details of our training and optimization procedure are provided in (Sec.
-3.2.3).
-3.1 Synthetic Dataset using Gaussian Avatar
-Figure 2: Visualization of the articulated mouse mesh and the learned geodesic embedding.
-Left: our mesh model consists of 140 articulated joints and 14,522 vertices, forming a kinematically consistent 3D structure for pose fitting.
-Right: color-coded visualization of the learned geodesic feature embedding, where principal components are mapped to HSV color space to enhance surface separability and promote reconstruction consistency.
-3.1.1 Mouse Parametric Model
-As shown in Fig 2, we employ the established mouse mesh model used in  [2] as avatar control signals.
-The model consists of $K=140$ articulated joints and 14,522 vertices, and conforms to Linear Blend Skinning (LBS) for pose control.
-The whole parameter set is denoted as $\Psi$, including joint rotations $\theta\in\mathbb{R}^{3K}$, global translation $t\in\mathbb{R}^{3}$, global rotation $r\in\mathbb{R}^{3}$, global scale $s\in\mathbb{R}$ and bone length deformation parameters $B\in\mathbb{R}^{21}$ (Please refer to  [2] for more details).
+Our goal is to reconstruct dense surface geometry of a laboratory mouse from a single image in a feedforward manner.
+We focus specifically on the C57BL/6 strain, a widely used model organism in biomedical research.
+We first introduce our Gaussian mouse avatar (AGAM) and synthetic dataset in Sec.
+1, then the structure of MoReMouse in Sec.
+2.
+1 Synthetic Dataset using Gaussian Avatar
+Figure 2: Left: the articulated mouse mesh we used. Right: color-coded geodesic feature embedding, promoting surface separability and reconstruction consistency.
+Mouse Parametric Model
+As shown in Fig 2, we modify the mouse model from  (49_bolanos2021three) following the approach of  (an2023three), where redundant mesh vertices (teeth and tone) are removed.
+The model consists of $N_{J}=140$ articulated joints and $N_{V}=13059$ vertices, and conforms to Linear Blend Skinning (LBS) for posing.
+Poses of this refined model serve as the avatar control signals.
+The whole parameter set is denoted as $\Psi$, including joint rotations $\theta\in\mathbb{R}^{3N_{J}}$, global translation $t\in\mathbb{R}^{3}$, global rotation $r\in\mathbb{R}^{3}$, global scale $s\in\mathbb{R}$ and bone length deformation parameters $B\in\mathbb{R}^{21}$ (Please refer to  (an2023three) for more details).
 We decompose $\Psi$ into local parameters $\Psi_{l}=\{\theta,B\}$ and global parameters $\Psi_{g}=\{r,t,s\}$ for the sake of simplicity.
-3.1.2 Gaussian Mouse Avatar
-Figure 3: Pipeline of the Gaussian Mouse Avatar Generation. The process begins with Mesh Fitting, followed by Gaussian model training.
-Gaussian point clouds [16] use an explicit parameterization for 3D representation.
+Gaussian Mouse Avatar
+Figure 3: Pipeline of the Gaussian Mouse Avatar Generation. The process begins with Mesh Fitting, followed by Gaussian Avatar training. “CNN” here is StyleUNet.
+Gaussian point clouds (kerbl2023gaussian) use an explicit parameterization for 3D representation.
 Each Gaussian point is characterized by a mean position $\mu\in\mathbb{R}^{3}$ and an anisotropic covariance matrix $\Sigma$:
@@
 During rendering, Gaussians within the pixel projection region are sorted by depth and composited using alpha blending.
-Inspired by Animatable Gaussians [20], we employ texture-based deformation transfer to train a neural representation of the Gaussian point cloud avatar, incorporating bias correction beyond the skeletal model, as shown in Fig.
+The pipeline of our AGAM is shown in Fig.
 3.
 Our key idea is to parameterize the mouse UV texture, mapping each valid texel in the UV space to a corresponding Gaussian point, where the texture color represents the Gaussian’s position.
-To obtain the canonical space representation, we set $\Psi_{g}=0$ and drive the mesh model using only $\Psi_{l}$, resulting in canonical vertex coordinates $V_{\Psi_{l}}$.
-The UV-mapped positions of these vertices are used to interpolate Gaussian locations in the UV texture via barycentric coordinates.
+To achieve global-pose invariant controlling of AGAM, we set $\Psi_{g}=0$ and drive the mesh model using only $\Psi_{l}$, resulting in canonical vertex coordinates $V_{\Psi_{l}}$.
+The UV-mapped positions of these vertices are used to interpolate Gaussian locations via barycentric coordinates.
 This defines the baseline Gaussian position map $\mu_{\Psi_{l}}$:
@@
 where $\mu_{0}$ is the UV coordinate map in the neutral pose, and $\mathcal{D}_{\Psi_{l}}$ represents the LBS deformation mapping induced by the skeletal model.
-During Gaussian optimization, we incorporate a CNN-based offset prediction module.
-A convolutional neural network StyleUNet [40] predicts the position offset map $\Delta\mu_{\Psi_{l}}$, conditioned on the UV coordinate map $\mu_{\Psi_{l}}$, to refine the Gaussian positions:
-$\mu^{\prime}_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}+\Delta\mu_{\Psi_{l}}).$ (3)
-The CNN also estimates additional Gaussian attributes, including color $c_{\Psi_{l}}$, opacity $o_{\Psi_{l}}$, covariance rotation parameters $q_{\Psi_{l}}$, and scaling parameters $s_{\Psi_{l}}$.
-Then the Gaussian Splatting rendering function generates the view-dependent rendered image $I_{c}^{\text{rend}}$:
+The core network of AGAM is a set of StyleUNets (wang2023styleavatar) that predict the position offset map $\Delta\mu_{\Psi_{l}}$ and other Gaussian attributes $c_{\Psi_{l}}$, $o_{\Psi_{l}}$, $q_{\Psi_{l}}$, and $s_{\Psi_{l}}$ conditioned on the UV coordinate map $\mu_{\Psi_{l}}$.
+Note that, the Gaussian positions $\mu^{\prime}_{\Psi_{l}}$ are derived by $\mu^{\prime}_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}+\Delta\mu_{\Psi_{l}}).$ Then the Gaussian Splatting rendering function generates the view-dependent rendered image $I_{c}^{\text{rend}}$:
 $I_{c}^{\text{rend}}=GS(R_{c},T_{c},K_{c},\mu^{\prime}_{\Psi_{l}},c_{\Psi_{l}},q^{\prime}_{\Psi_{l}},s_{\Psi_{l}}),$ (3)
 where $GS(\cdot)$ denotes the Gaussian splatting renderer, and $R_{c},T_{c},K_{c}$ are camera parameters.
-3.1.3 Avatar Training
-To train the Gaussian avatar, we define loss functions between the real image $I_{c}$ and the rendered image $I_{c}^{\text{rend}}$:
+Avatar Training
+We define loss functions between the real image $I_{c}$ and the rendered image $I_{c}^{\text{rend}}$ as:
 $\mathcal{L}$ $=\mathcal{L}_{1}(I_{c}-I_{c}^{\text{rend}})+\lambda_{\text{SSIM}}(1-\mathcal{L}_{\text{SSIM}}(I_{c},I_{c}^{\text{rend}}))$ (4)
 $\quad+\lambda_{\text{LPIPS}}\mathcal{L}_{\text{LPIPS}}(I_{c},I_{c}^{\text{rend}}),$
-where $\mathcal{L}_{\text{SSIM}}$ quantifies structural similarity, while $\mathcal{L}_{\text{LPIPS}}$ assesses perceptual similarity.
-The loss weights are set to $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{LPIPS}}=0.1$.
-Additionally, we enforce Total Variation (TV) loss to regularize Gaussian point cloud smoothness.
-We use a 6-view mouse video sequence ”markerless_mouse_1” proposed in [11] for avatar training.
-Each sequence captures 18,000 frames of a single freely moving mouse in an open field at 1152x1024@100FPS, where the mouse mesh models are provided by  [2].
-Note that, we only use 800 frames for training, which are uniformly sampled from the first 8000 frames of ”markerless_mouse_1”.
+where $\mathcal{L}_{\text{SSIM}}$ is structural similarity loss, $\mathcal{L}_{\text{LPIPS}}$ is perceptual similarity loss, $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{LPIPS}}=0.1$.
+We use Total Variation (TV) loss to enforce point smoothness.
+See the supplementary material for further details.
+Since C57BL/6 mice usually show similar texture, we only use a 6-view mouse video sequence “markerless_mouse_1” proposed by (dunn2021geometric) for avatar training.
+This sequence captures 18,000 frames of a single freely moving mouse in an open field at 1152x1024@100FPS, where the mouse mesh models are borrowed from  (an2023three).
+Note that, we only use 800 frames for training, which are uniformly sampled from the first 8000 frames of “markerless_mouse_1”.
 Avatar training takes 400k steps.
-3.2 MoReMouse Architecture
-3.2.1 Transformer and Triplane Representation
-Similar to LRM [14], MoReMouse adopts a Transformer and triplane-based architecture, designed for monocular 3D reconstruction.
-The model consists of three key components: an image encoder, an image-to-triplane decoder, and a triplane-based geometric representation that integrates both NeRF [25] and DMTet [35] for implicit and explicit geometry modeling.
-First, the image encoder employs a pre-trained DINOv1 backbone to extract a latent feature vector from the input image.
-Then, a large Transformer decoder employs cross-attention to project these image features into a 3D triplane representation.
-This triplane-based geometric representation encodes color, density, and additional spatial attributes at any queried 3D location.
-During the initial training phase, to regularize the convergence of the spatial density field, we do not introduce explicit geometry.
-Instead, we employ NeRF for differentiable volumetric rendering, allowing gradients to propagate throughout the entire space.
-Once the density field has sufficiently converged, we introduce DMTet as an explicit geometric representation for further refinement.
-DMTet provides well-defined surfaces and enables direct geometric supervision.
+Dense-view Dataset
+We drive AGAM using the poses from all frames of the “markerless_mouse_1” video sequence, and render evenly distributed 64-view synthetic images.
+Among the full 18,000 frames, we use the first 6,000 frames for training and the last 6,000 frames for testing.
+Note that, the 800 frames used for avatar creation all belong to the first 6000 frames, therefore the poses in testset are unseen.
+To construct the training set, we sample two sets of 64 viewpoints per frame, rendering 12,000 multi-view scenes in total.
+The optical axes of these viewpoints point at the scene origin.
+Please refer to the supplementary file for more information about view selection and data samples.
+2 MoReMouse Architecture
+Transformer and Triplane Representation
+Similar to LRM (hong2023lrm), MoReMouse adopts a Transformer and triplane-based architecture, designed for monocular 3D reconstruction.
+As illustrated in Fig.
+4, the model consists of three key components: an image encoder, an image-to-triplane decoder, and a triplane-based geometric representation that integrates both NeRF (mildenhall2021nerf) and DMTet (shen2021deep) for implicit and explicit geometry modeling.
+We use a pretrained DINOv2 encoder to extract latent features from the input image.
+These features are decoded via a transformer with cross-attention to generate a triplane representation, which encodes color, density, and semantic attributes for any 3D query point.
+During the initial training stage, we employ NeRF for volumetric rendering without introducing explicit geometry, enabling smooth gradient propagation through the entire density field.
+After convergence, we switch to DMTet to obtain an explicit surface representation for fine-level geometric supervision.
 The fine-scale structures of mice, particularly in the limbs and tail, present challenges for high-fidelity reconstruction in a feed-forward manner.
@@
 Additional architectural details are provided in the supplementary materials.
-Figure 4: Architecture of the MoReMouse reconstruction network.
-Given a monocular input image, a DINO-V2 encoder extracts high-level image features, which are processed by a transformer-based decoder to generate a triplane representation.
-These triplanes, combined with learnable positional embeddings, are queried by a Multi-Head MLP to predict per-point color, density, and geodesic embedding.
-Rendering is performed via either NeRF (for volumetric supervision) or DMTet (for explicit geometry output), producing RGB images, color-coded embeddings, and surface meshes.
-3.2.2 Feature Embedding for Surface Consistency
-The reconstruction of mice poses significant challenges due to their black fur, which lacks distinctive surface features, and the presence of intricate fine-scale geometric details.
-Inspired by Animatable Gaussians [20], we incorporate feature embeddings into the mouse surface representation to enhance reconstruction stability.
-These embeddings serve as additional color channels during training, acting as supervisory signals for MoReMouse rendering.
-Specifically, given the set of mouse model vertices $V\in\mathbb{R}^{13059\times 3}$, we compute the pairwise geodesic distances between vertices to obtain the geodesic distance matrix:
-$G_{V}\in\mathbb{R}^{13059\times 13059},\quad G_{V}(i,j)=d_{\text{geo}}(V_{i},V_{j}),$ (6)
+Figure 4: Overview of the MoReMouse architecture.
+Given a single image, a DINOv2 encoder and transformer decoder generate triplane features, which are queried to produce color, density, and embeddings.
+Rendering is performed via NeRF or DMTet for volumetric or surface outputs.
+Feature Embedding for Surface Consistency
+Mouse reconstruction is challenging due to black fur with weak texture cues and fine-scale surface details.
+Inspired by Animatable Gaussians (li2024animatable), we incorporate feature embeddings computed from geodesic distances (shamai2017geodesic) into the surface representation.
+These embeddings are used as additional color channels during training, serving as semantic priors to enhance reconstruction stability.
+As shown in Fig. 2, given the set of mouse model vertices $V\in\mathbb{R}^{N_{V}\times 3}$, we compute the pairwise geodesic distances between vertices to obtain the geodesic distance matrix:
+$G_{V}\in\mathbb{R}^{N_{V}\times N_{V}},\quad G_{V}(i,j)=d_{\text{geo}}(V_{i},V_{j}),$ (5)
 where $d_{\text{geo}}(V_{i},V_{j})$ represents the geodesic distance between vertices $V_{i}$ and $V_{j}$.
-To learn a continuous correspondence embedding that captures the intrinsic structure of the mouse surface, we define the embedding matrix $E\in\mathbb{R}^{13059\times 3}$ and optimize it by minimizing the $L_{2}$ loss between the embedding-derived distance matrix $G_{E}$ and the geodesic distance matrix $G_{V}$:
+To learn a continuous correspondence embedding that captures the intrinsic structure of the mouse surface, we define the embedding matrix $E\in\mathbb{R}^{N_{V}\times 3}$ and optimize it by minimizing the $L_{2}$ loss between the embedding-derived distance matrix $G_{E}$ and the geodesic distance matrix $G_{V}$:
 $\mathcal{L}_{\text{geo}}=\sum_{i,j}\left\|G_{E}(i,j)-G_{V}(i,j)\right\|_{2}^{2},$ (6)
@@
 This transformation improves the separability of the embeddings in the rendered images.
-Notably, a simple rotation of the embeddings in color space can generate new feature mappings:
-$E^{\prime}=R(\theta)E,$ (9)
-where $R(\theta)$ is a rotation matrix.
-Extending the embeddings to higher dimensions could further improve reconstruction robustness.
-However, we leave the exploration of such extensions to future work.
-3.2.3 Training and Optimization
-Following the method outlined in Section 3.1.2, we generate freely renderable Mouse Avatars for each frame of the ”markerless_mouse_1” video sequence.
-As the dataset consists of 18,000 frames, we use the first 6,000 frames for training and the last 6,000 frames for testing.
-To construct the training set, we sample two sets of 64 viewpoints per frame, rendering 12,000 multi-view scenes in total.
-The optical axes of these viewpoints converge at the scene origin.
+Training and Optimization
 During training, two randomly selected viewpoints from each scene serve as the input-output pair.
@@
 We employ a two-stage training strategy, where NeRF-based volumetric rendering is first trained for 60 epochs to ensure proper density field convergence, followed by DMTet-based explicit geometry modeling for an additional 100 epochs to refine surface details.
-The total training loss is defined as:
-$\mathcal{L}$ $=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{smoothL1}}\mathcal{L}_{\text{smoothL1}}$ (10)
-$\quad+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{LPIPS}}\mathcal{L}_{\text{LPIPS}}.$
-Here, $\mathcal{L}_{\text{MSE}}$ enforces pixel-wise supervision for both RGB values and feature embeddings, while $\mathcal{L}_{\text{depth}}$ ensures depth consistency within the valid object region.
-The smooth L1 loss $\mathcal{L}_{\text{smoothL1}}$ penalizes large RGB discrepancies, and $\mathcal{L}_{\text{mask}}$ applies binary cross-entropy supervision to the predicted opacity maps.
-The perceptual loss $\mathcal{L}_{\text{LPIPS}}$ measures high-level perceptual similarity in a deep feature space.
-The hyperparameters $\lambda$ control the relative importance of each term.
+Details of the training loss terms and hyperparameters are provided in the supplementary material.
 4 Experiments
-4.1 Experimental Setup
+1 Experimental Setup
 We evaluate MoReMouse on two types of testing datasets: a synthetic benchmark and a real-world captured dataset.
 Synthetic Evaluation Data.
-We use the last 6,000 frames from the markerless_mouse_1 sequence to generate a synthetic multi-view dataset consisting of 4 orthogonal viewpoints.
+We use the last 6,000 frames from the “markerless_mouse_1” sequence to generate a synthetic multi-view dataset consisting of 4 orthogonal viewpoints.
 These views are rendered using our Gaussian avatar pipeline and serve as a controlled evaluation set where ground-truth camera parameters and rendering conditions are known.
 Real Captured Data.
-To assess performance under realistic conditions, we collect a real-world multi-view dataset using industrial cameras.
+To assess performance under realistic conditions, we collect a real-world multi-view dataset using PointGray industrial cameras.
 This dataset contains 5,400 frames of freely moving mice captured from four calibrated viewpoints.
-The camera rotation angles relative to the reference view are $0^{\circ}$, $-33.39^{\circ}$, $-78.60^{\circ}$, and $-181.02^{\circ}$, respectively.
+More details about camera setup and calibration are provided in the supplementary materials.
 Baselines.
 Due to the lack of monocular 3D reconstruction methods specifically designed for mice, we compare MoReMouse against four general-purpose single-view reconstruction baselines:
-  • TripoSR [39]: A state-of-the-art open-source Large Reconstruction Model (LRM) known for strong single-view 3D generation performance.
-  • Triplane-GS [50]: A LRM variant based on Gaussian representations.
-  • LGM [38]: A multi-view diffusion-based method that reconstructs 3D Gaussian point clouds.
-  • InstantMesh [45]: A mesh-based 3D reconstruction method using multi-view diffusion priors.
-For synthetic data, we additionally evaluate the multi-view diffusion baselines (LGM and InstantMesh) using ground-truth views as input, which serves as an approximate upper bound for multi-view reconstruction.
-Evaluation Protocol.
+  • TripoSR (tochilkin2024triposr): A state-of-the-art open-source Large Reconstruction Model (LRM) known for strong single-view 3D generation performance.
+  • Triplane-GS (zou2024triplane): A LRM variant based on Gaussian representations.
+  • LGM (tang2024lgm): A multi-view diffusion-based method that reconstructs 3D Gaussian point clouds.
+  • InstantMesh (xu2024instantmesh): A mesh-based 3D reconstruction method using multi-view diffusion priors.
 All methods are evaluated on the task of novel view synthesis, since obtaining ground-truth 3D geometry for mouse subjects is impractical.
-For each frame, a single view is used as input, and the remaining three views are used to evaluate the visual fidelity of synthesized images.
-We report three standard image-based metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS).
-4.2 Experimental Results
+Additionally, we include surface geometry metrics (IoU against multi-view visual hulls) in the supplementary material.
+We use Track-Anything (yang2023trackanythingsegmentmeets) to remove background regions in real videos.
+Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+[…]
+TripoSR 13.673 0.8032 0.18255
+[…]
+LGM 14.460 0.8805 0.13274
+[…]
+[…]
+TripoSR 11.518 0.8114 0.19672
+[…]
+[…]
+[…]
+Table 1: Quantitative results on both synthetic and real-world datasets. MoReMouse achieves the best performance across all metrics in both domains.
+Figure 5: Qualitative comparison of novel view synthesis results on synthetic data.
+MoReMouse produces coherent and anatomically plausible reconstructions, accurately capturing mouse posture and geometry, while baseline methods often exhibit structural distortions.
+Figure 6: Qualitative comparison on real data.
+MoReMouse maintains consistent geometry and appearance, while baselines often show pose and shape artifacts.
+Occasional tail discontinuities arise from imperfect Track-Anything masks under reflections or occlusions; our model remains robust and recovers plausible tail shapes.
+2 Experimental Results
 Quantitative Results.
@@
 For each metric, we highlight the top-3 performing methods using color shading, with darker tones indicating better performance.
-Table 1: Quantitative results on both synthetic and real-world datasets.
-MoReMouse achieves the best performance across all metrics in both domains, significantly outperforming baseline methods in terms of reconstruction fidelity.
-Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-synthetic Ours 22.027 0.9660 0.05279
-TripoSR 13.673 0.8032 0.18255
-Triplane-GS 18.049 0.9268 0.10151
-LGM 14.460 0.8805 0.13274
-InstantMesh 15.821 0.8987 0.11312
-real Ours 18.422 0.9478 0.08674
-TripoSR 11.518 0.8114 0.19672
-Triplane-GS 16.789 0.9298 0.11002
-LGM 15.215 0.9197 0.12838
-InstantMesh 15.631 0.9175 0.11339
 Across all three metrics for novel view synthesis, MoReMouse consistently outperforms the baseline methods, demonstrating its ability to effectively infer plausible mouse geometry and pose from a single image.
-Notably, in Table 2, MoReMouse uses only a single input view, while LGM and InstantMesh are evaluated with 4 and 6 views respectively.
-For LGM, the 4 input views happen to coincide with the 4 evaluation viewpoints, offering an idealized setting that may favor the method.
-Nonetheless, MoReMouse achieves comparable performance despite relying on significantly less input information, underscoring its strong generalization capabilities in monocular reconstruction scenarios.
 We also observe that the quantitative metrics on the real-world dataset are generally lower than on the synthetic dataset, especially for PSNR.
@@
 This confirms that models trained on synthetic data can generalize to real-world mouse motion capture scenarios.
-Table 2: Comparison with multi-view diffusion-based methods using ground-truth viewpoints on the synthetic dataset.
-Despite using only a single view, MoReMouse performs competitively with multi-view approaches that use 4 or 6 calibrated inputs.
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Ours (single view) 22.027 0.9660 0.05279
-LGM (4 views) 23.998 0.9698 0.06458
-InstantMesh (6 views) 28.295 0.9831 0.02716
+Full tables of IoU metrics, expanded baseline comparisons, and runtime/memory usage are provided in the supplementary.
 Qualitative Results.
@@
 In contrast, baseline methods often exhibit structural distortions in novel views and inconsistent pose transitions across frames.
-Diffusion-based approaches such as InstantMesh and LGM tend to produce view-inconsistent geometry, likely due to the absence of strong priors on mouse anatomy.
-Single-view approaches like Triplane-GS and TripoSR frequently fail to reconstruct accurate poses, especially under extreme articulations or occlusions.
-4.3 Ablation Study
-4.3.1 Comparison with Finetuned TripoSR
-In previous experiments, we compared MoReMouse with four general-purpose single-view 3D reconstruction baselines, demonstrating the effectiveness of our complete training pipeline.
-However, these reconstruction models were pre-trained on datasets that may not contain sufficient small animal or mouse-like samples, which could limit their performance on our mouse-specific benchmarks.
-To assess whether the improvements of MoReMouse are due to architectural and representation innovations rather than only data augmentation, we fine-tune TripoSR on our mouse dataset using a similar training protocol.
-Specifically, we train the model for 120 epochs with a learning rate of $1\times 10^{-6}$.
-As shown in Table 3 and illustrated in Fig.
-7, the fine-tuned TripoSR achieves improved performance compared to its original version, but still lags behind MoReMouse across most metrics.
-This suggests that our model architecture, including the triplane encoding and geodesic feature embedding, plays a critical role in achieving high-fidelity reconstructions.
+3 Ablation Study
+Comparison with Finetuned TripoSR
+To further validate the benefits of our architectural design, we fine-tune TripoSR on our mouse dataset for 120 epochs with a learning rate of $1\times 10^{-6}$.
+As shown in Table 2 and Fig.
+7, the fine-tuned TripoSR shows noticeable improvement over its original version but still underperforms MoReMouse across most metrics.
+This indicates that, beyond data adaptation, our architectural choices—such as triplane encoding and geodesic embeddings—contribute significantly to reconstruction quality.
+We also fine-tune Triplane-GS and report its performance in the supplementary, which reveals reconstruction artifacts particularly around limbs and the tail.
+Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+[…]
+[…]
+[…]
+TripoSR-Tuned 18.162 0.9461 0.09354
 Table 2: Comparison between MoReMouse and TripoSR finetuned on mouse data.
-Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-synthetic Ours 22.027 0.9660 0.05279
-TripoSR-Tuned 21.996 0.9676 0.06333
-real Ours 18.422 0.9478 0.08674
-TripoSR-Tuned 18.162 0.9461 0.09354
-4.3.2 Effect of Feature Embedding
+Effect of Feature Embedding
 We evaluate the effectiveness of the geodesic feature embedding module by training a variant of MoReMouse without this component.
@@
 The geodesic embedding encourages coherent geometry and finer detail recovery.
-Figure 5: Qualitative comparison of novel view synthesis results on synthetic data.
-MoReMouse produces coherent and anatomically plausible reconstructions, accurately capturing mouse posture and geometry, while baseline methods often exhibit structural distortions.
-Figure 6: Qualitative comparison of novel view synthesis results on real data.
-MoReMouse generalizes well to real-world mouse recordings, maintaining consistent geometry and appearance across views, while baseline methods often suffer from inconsistent shapes and misaligned postures.
 Figure 7: Qualitative comparison between our MoReMouse method and a finetuned TripoSR baseline.
@@
 Embeddings help guide fine-grained surface details and improve anatomical accuracy.
+Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+[…]
+w/o embedding 21.805 0.9655 0.05467
+[…]
+w/o embedding 18.250 0.9469 0.08767
 Table 3: Ablation study on the effect of feature embedding. Removing embedding results in lower reconstruction quality, especially in perceptual similarity.
-Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-synthetic full model 22.027 0.9660 0.05279
-w/o embedding 21.805 0.9655 0.05467
-real full model 18.422 0.9478 0.08674
-w/o embedding 18.250 0.9469 0.08767
-4.4 Failure cases
-Figure 9 showcases typical failure cases on real-world data.
-The first row highlights the model’s inability to generalize to rare upright postures, due to the absence of such poses in the training set.
-This leads to implausible geometry and emphasizes the need for more pose-diverse avatar data.
-In the second row, reconstruction fails when the tail occludes the head—especially under shaved-fur conditions—causing the model to misinterpret head structure.
-While our preprocessing removes background, we do not explicitly account for occlusions from scene objects or self-occlusion, which can degrade reconstruction fidelity.
-Additionally, we observe that small regions such as feet and snouts are more prone to distortion, likely due to limited resolution in those areas.
-Future work may explore resolution-aware refinement or occlusion-robust modeling to address these limitations.
-Figure 9: Representative failure cases.
-Top: failure due to lack of standing mouse data in training, leading to unrealistic reconstructions.
-Bottom: failure due to tail-induced occlusion from shaved fur, which interferes with head feature extraction.
 5 Discussion
@@
 Ablation studies further validate the architectural contributions of our model beyond dataset-specific priors.
-We hope that MoReMouse can serve as a foundation for future research in computational ethology and non-rigid animal modeling, promoting more precise, efficient, and scalable behavioral analysis tools for biomedical research.
+We hope that MoReMouse can serve as a foundation for future research in computational ethology and animal modeling, promoting more precise, efficient, and scalable behavioral analysis tools for biomedical research.
 Limitations.
@@
 This design limits our ability to track full 3D poses in a real-world coordinate system.
-Ethical Statement.
+Ethical Statement
+Please refer to supplementary file for Ethical Statements.
+Acknowledgements
+This work was supported by the National Natural Science Foundation of China (No. 62125107) and the Shuimu Tsinghua Scholar Program (2024SM324).
+Appendix
+Figure A1: Pipeline for constructing the Dense-view Dataset and synthetic evaluation data.
+The first 6,000 frames of the “markerless_mouse_1” 6-view sequence are used to build the dense-view dataset for training, while the last 6,000 frames are reserved for generating the synthetic evaluation set.
+For each training frame, AGAM is rendered from 64 randomly sampled viewpoints on a sphere of radius $2.22$, resulting in 12,00012{,}000 multi-view scenes.
+Appendix A1 Ethical Statement
 All procedures involving mice in this study were non-invasive and observational, involving only video recording of natural behaviors without any physical intervention or stress to the animals.
@@
 The study adhered to the 3Rs principle (Replacement, Reduction, Refinement) in animal research.
-Acknowledgements
-This work was supported by the National Natural Science Foundation of China (No. 62125107) and the Shuimu Tsinghua Scholar Program (2024SM324).
-Supplementary Material
-1 MoReMouse Model Details
-We summarize the architectural and training hyperparameters of the MoReMouse model in Table 5, covering key components such as the image encoder, triplane representation modules, multi-head MLP, NeRF and DMTet renderers, and training configurations.
-Table 5: MoReMouse Architecture and Training Hyperparameters
+Appendix A2 Dataset Construction
+1 Dense-view Dataset Generation
+As illustrated in Fig.
+A1, we construct our Dense-view Dataset using the first 6,000 frames of the “markerless_mouse_1” video sequence.
+For each frame, we sample two independent sets of $64$ viewpoints, producing 12,00012{,}000 multi-view scenes in total, each containing $64$ images.
+For viewpoint sampling, we first align the local coordinate origin of the generated Animatable Gaussian Avatar of Mouse (AGAM)—given by the origin of its internally driven mesh model—to the world coordinate origin.
+The Gaussian mouse model is uniformly scaled down by a factor of $180$ so that the mouse fits inside a unit sphere.
+We then randomly sample camera centers on a sphere of radius $2.22$, allowing for both top-down and bottom-up observations, so our dataset can accommodate to scenarios where mice are recorded from above or through a transparent floor.
+The rendered images have a resolution of $800\times 800$ pixels, and the FoV is fixed at $29.86^{\circ}$.
+2 Synthetic Evaluation Dataset Generation
+We use the last 6,000 frames from the “markerless_mouse_1” sequence to construct a Synthetic Evaluation Dataset consisting of four orthogonal viewpoints (front, back, left, and right).
+Following a similar placement strategy as in the Dense-view Dataset, we align the local coordinate origin of the Gaussian mouse model with the world coordinate origin and uniformly scale it down by a factor of $180$ so that it fits within a unit sphere.
+The four cameras are positioned on a sphere of radius $2.22$ with optical axes pointing toward the scene origin, evenly spaced in azimuth by $90^{\circ}$.
+The rendered images use a fixed FoV of $29.86^{\circ}$ and have a resolution of $800\times 800$.
+To verify that the poses in the evaluation set differ from those in the training set, we visualize the pose distributions of the Dense-view Dataset and Synthetic Evaluation Dataset using t-SNE (maaten2008visualizing), as shown in Fig.
+A2.
+Note that the evaluation set contains a large number of unseen poses, ensuring that the reported performance reflects the model’s ability to generalize to novel poses.
+Figure A2: t-SNE visualization of mouse poses.
+The visualization shows that the training and testing poses are separated in the embedding space.
+Blue points are poses from Dense-view training dataset, red points are poses from the evaluation set.
+3 Real Captured Data
+As shown in Fig.
+A3, we collect a real-world multi-view dataset using PointGrey industrial cameras.
+The capture setup consists of four synchronized cameras positioned around a freely moving C57BL/6 mouse.
+The mouse is placed inside a transparent cylinder with a diameter of 20cm and a height of 20cm.
+Each camera records at a resolution of $2048\times 1500$ pixels, with a horizontal field-of-view (FoV) of $22.34^{\circ}$ and a vertical FoV of $16.44^{\circ}$.
+This real-world dataset serves to evaluate the generalization of our model from synthetic training data to realistic environments.
+Before feeding the images to the network, all mice are cropped out with background removed.
+Figure A3: Real-world dataset acquisition setup.
+Four PointGrey industrial cameras are arranged to capture a freely moving C57BL/6 mice.
+This dataset enables the evaluation of cross-domain generalization from synthetic to real-world conditions.
+Figure A4: Qualitative comparison on the original 6-view “markerless_mouse_1” dataset.
+Our method produces more geometrically accurate and visually consistent reconstructions compared to other baselines.
+We also observe that TRELLIS occasionally outputs nearly blank images, indicating instability when handling mouse poses.
+Figure A5: Qualitative comparison on the real-world multi-view dataset.
+Our method reconstructs anatomically coherent geometry and consistent appearance across views, whereas competing methods often produce structural distortions or inconsistent textures.
+Appendix A3 More Experimental Results
+In this section, we present additional experimental results beyond those reported in the main paper.
+In addition to the four baseline methods compared in the main text (TripoSR (tochilkin2024triposr), Triplane-GS (zou2024triplane), LGM (tang2024lgm), and InstantMesh (xu2024instantmesh)), we further compare our method against two recent reconstruction approaches:
+  • 3D Fauna (li2024learning): A template-free method for learning 3D reconstruction of animals from online monocular images.
+  • TRELLIS (xiang2025structured): A structured generative model designed to capture fine-grained articulations through a hierarchical representation.
+We also report a surface geometry metric based on the IoU between the rendered reconstruction and multi-view visual-hull masks, offering a direct measure of geometric alignment.
+To further examine the behavior of Triplane-GS on our task, we fine-tune it on our mouse dataset for 200 epochs using a learning rate of $1\times 10^{-6}$.
+However, the resulting Gaussian point clouds remain blurry and lack sharp geometric boundaries, making it difficult to capture fine structures such as limbs and tails.
+We additionally adapt 3D-Fauna’s pan-category deformable model to our 6-view mouse video dataset via fine-tuning.
+During training, each iteration samples data with a probability of $0.6$ from the original 3D-Fauna dataset and $0.4$ from our mouse dataset.
+We freeze the base shape bank while fine-tuning the instance deformation network, pose predictor, and texture field.
+Optimization is performed for 500k iterations with a batch size of 10, using Adam (learning rate $1\times 10^{-4}$ for the instance network, $1\times 10^{-3}$ for the base network, weight decay 0).
+However, we observe that 3D-Fauna struggles to fit our mouse dataset, which covers the entire viewing sphere, leading to suboptimal performance in multi-view reconstruction.
+Finally, we benchmark inference efficiency: MoReMouse runs at approximately 0.91 seconds per frame on an NVIDIA A100 GPU, consuming 7.9 GB of GPU memory during inference.
+1 Comparison on Original Images of “markerless_mouse_1”
+In addition to the evaluations on the Synthetic Evaluation Data constructed from the last 6,000 frames of the “markerless_mouse_1” sequence, we further evaluate all methods by directly reconstructing the same 6,000 frames from the original 6-view captures of “markerless_mouse_1”.
+Qualitative comparisons are shown in Fig.
+A4, and quantitative results are summarized in Table A1.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
+[…]
+[…]
+TripoSR 15.837 0.9202 0.1453 0.4474
+[…]
+Triplane-GS 15.703 0.9182 0.1398 0.4682
+Triplane-GS-Tuned 17.549 0.9293 0.1117 0.5966
+LGM 14.079 0.9051 0.1519 0.3230
+InstantMesh 14.876 0.9127 0.1313 0.4080
+Fauna 14.261 0.9086 0.1488 0.4030
+Fauna-Tuned 14.054 0.9089 0.1481 0.2956
+TRELLIS 13.199 0.9146 0.1304 0.1242
+Table A1: Quantitative results on the original 6-view “markerless_mouse_1” dataset. Cyan shading indicates top-3 performance.
+2 Comparison on Real Data
+we evaluate all compared approaches on our captured real multi-view dataset.
+Qualitative comparisons are presented in Figure A5, and quantitative results are summarized in Table A2.
+Our method consistently achieves the best performance across all three metrics, indicating superior reconstruction fidelity and perceptual quality.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
+[…]
+[…]
+TripoSR 11.518 0.8114 0.19672 0.4741
+[…]
+Triplane-GS 16.789 0.9298 0.11002 0.5743
+Triplane-GS-Tuned 18.628 0.9469 0.09239 0.6099
+LGM 15.215 0.9197 0.12838 0.4138
+InstantMesh 15.631 0.9175 0.11339 0.5466
+3D Fauna 13.284 0.8837 0.12890 0.4467
+3D Fauna-Tuned 13.726 0.8977 0.12038 0.4254
+TRELLIS 15.087 0.9279 0.10968 0.3095
+Table A2: Quantitative results on the real-world multi-view dataset. Cyan shading indicates top-3 performance.
+Appendix A4 Failure cases
+Figure A6 showcases typical failure cases on real-world data.
+The first row highlights the model’s inability to generalize to rare upright postures, due to the absence of such poses in the training set.
+This leads to implausible geometry and emphasizes the need for more pose-diverse avatar data.
+In the second row, reconstruction fails when the tail occludes the head—especially under shaved-fur conditions—causing the model to misinterpret head structure.
+While our preprocessing removes background, we do not explicitly account for occlusions from scene objects or self-occlusion, which can degrade reconstruction fidelity.
+Additionally, we observe that small regions such as feet and snouts are more prone to distortion, likely due to limited resolution in those areas.
+Future work may explore resolution-aware refinement or occlusion-robust modeling to address these limitations.
+Figure A6: Representative failure cases.
+Top: failure due to lack of standing mouse data in training, leading to unrealistic reconstructions.
+Bottom: failure due to tail-induced occlusion from shaved fur, which interferes with head feature extraction.
+Appendix A5 More Details about MoReMouse Training
+Our model is trained in two stages: the NeRF-based volumetric rendering stage is trained for $60$ epochs to ensure proper density field convergence, followed by the DMTet-based explicit geometry stage trained for $100$ epochs to refine surface details.
+The overall optimization objective combines geometric, photometric, and perceptual terms, and is defined as:
+$\mathcal{L}$ $=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{smoothL1}}\mathcal{L}_{\text{smoothL1}}$ (8)
+$\quad+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{LPIPS}}\mathcal{L}_{\text{LPIPS}}.$
+Here, $\mathcal{L}_{\text{MSE}}$ enforces pixel-wise supervision for both RGB values and feature embeddings, while $\mathcal{L}_{\text{depth}}$ ensures depth consistency within the valid object region.
+The smooth L1 loss $\mathcal{L}_{\text{smoothL1}}$ penalizes large RGB discrepancies, and $\mathcal{L}_{\text{mask}}$ applies binary cross-entropy supervision to the predicted opacity maps.
+The perceptual loss $\mathcal{L}_{\text{LPIPS}}$ measures high-level perceptual similarity in a deep feature space.
+The hyperparameters $\lambda$ control the relative importance of each term.
+We summarize the architectural and training hyperparameters of the MoReMouse model in Table A3, covering key components such as the image encoder, triplane representation modules, multi-head MLP, NeRF and DMTet renderers, and training configurations.
 Module Parameter Value
@@
 Shared hidden layers $10$
-Hidden layers for density/feature/deformation heads $1$
+Hidden layers for density head 11
+Hidden layers for feature/deformation heads 11
 Renderer-NeRF Samples per ray $128$
@@
 Density activation trunc_exp
-Density bias $-1.0$
-Renderer-DMTet SDF bias $-4.0$
+Density bias 1.0-1.0
+Renderer-DMTet SDF bias 4.0-4.0
 Isosurface resolution $256^{3}$
@@
 Warm-up steps $3000$
-$\lambda_{\text{mse}}$, $\lambda_{\text{LPIPS}}$ $1.0$
+$\lambda_{\text{mse}}$ $1.0$
+$\lambda_{\text{LPIPS}}$ $1.0$
 $\lambda_{\text{mask}}$ $0.3$
-$\lambda_{\text{smooth}}$, $\lambda_{\text{depth}}$ $0.2$
+$\lambda_{\text{smooth}}$ $0.2$
+$\lambda_{\text{depth}}$ $0.2$
+Table A3: MoReMouse Architecture and Training Hyperparameters